Westerville Company reported the following results from last year’s operations:

Sales $ 1,000,000
Variable expenses 300,000
Contribution margin 700,000
Fixed expenses 500,000
Net operating income $ 200,000
Average operating assets $ 625,000

At the beginning of this year, the company has a $120,000 investment opportunity with the following cost and revenue characteristics:

  

Sales $ 200,000
Contribution margin ratio 60 % of sales
Fixed expenses $ 90,000

The company’s minimum required rate of return is 15%.

Questions:

4. What is the margin related to this year’s investment opportunity?

7. If the company pursues the investment opportunity and otherwise performs the same as last year, what margin will it earn this year?

9. If the company pursues the investment opportunity and otherwise performs the same as last year, what ROI will it earn this year?

11. What is last year’s residual income?
Answer #1

Net income for investment = 200000*60%-90000 = 30000

4) Margin = 30000/200000 = 15%

7) Margin = 230000/1200000 = 19.17%

9) ROI = 230000/745000 = 30.87%

11) Residual income = 200000-(625000*15%) = 106250